# Scaled Block Vecchia Approximation for High-Dimensional Gaussian Process Emulation on GPUs

### v2

@@ 5.1.2. Random Anchor Clustering (RAC)
4:  Randomly choose $k_{p}$ local centers $C_{p}=\{\bm{c}_{p,i}\mid i=1,2,\dots,k_{p}\}$ in $X_{p}$, and initialize $B_{p,j}=\{\bm{c}_{p,i}\}.$
5:  Assign $\bm{x}_{p,i}$ to $j$th cluster $B_{p,j}$, where $j=\arg\min_{j}\|\bm{x}_{p,i}-\bm{c}_{p,j}\|^{2},j=1,2,\dots,k_{q}$.
@@
7:Return: Clusters $B_{1:P}$
5.1.3. Filtered $m$-Nearest Neighbor Search (NNS)
With block sets $\bm{B}_{1:P}$, the next goal is to find the exact $m$ nearest neighbors for all block centers in high-dimensional space, where search areas are varied for each query center in the Vecchia-based GPs, see Equation (2). Existing methods (Abbasifard et al., 2014; Ferhatosmanoglu et al., 2001; Clarkson et al., 2006) for exact $m$-NNS with query-specific search areas face limitations, especially for large datasets.
To address these issues, we adopt the filtered $m$-NNS on each block $B_{p,i}$, detailed in Algorithm 4. This approach selects a small filtered subset (a circle of radius $\lambda$) for brute-force $m$-NNS to obtain exact results while reducing computational cost. In Algorithms 1 and 4, $\mathcal{V}$ denotes the filtered $m$-NNS function, which begins by computing a Monte Carlo-based distance threshold. This threshold $\lambda$ is calculated using:
@@ 5.1.3. Filtered $m$-Nearest Neighbor Search (NNS)
Here, $n$ is the total number of points, $m$ is the number of nearest neighbors, $d$ is the dimensionality, and $\alpha$ is an expansion factor to account for irregularities in point distribution (e.g., $\alpha=100$ implies the candidate set is expected to be 100 times larger than $m$). Then, we prepare the small filtered subset, which consists of two steps, coarser and finer preparation. In the coarser preparation, each worker redistributes candidate clusters to the corresponding worker if the distance between two centers is within the calculated threshold using MPI_Alltoall. The complementary coarser candidate set is denoted as $B_{p}^{cand}$ (lines 10-13 in Algorithm 4). In the finer partition, each worker independently selects finer candidates $S_{p,i}$ for the $i$th cluster, i.e., $\bm{x}_{k}$ is added to the finer candidate set $S_{p,i}$ if the distance between the $i$th local center and the data point $\bm{x}_{k}$ is within the threshold $\lambda$ (lines 15-22 in Algorithm 4). Finally, each local center $\bm{c}_{i}$ applies brute-force search for its $m$ nearest neighbors $J_{p,i}$ from the finer candidates set $S_{p,i}$. The computational cost of brute search in $S_{p,i}$ is negligible compared to $\mathcal{O}(n)$ due to $\alpha m\ll n$ and total block count $K\ll n$, thus filtered $m$-NNS achieves computational complexity $\mathcal{O}(\alpha m+K)$ and has a simple implementation.
Figure 3 illustrates the filtered m-NNS using an example of four workers. In the coarser candidate preparation, MPI_Alltoall is utilized to distribute related blocks (within distance threshold $\lambda$) to the 33rd worker, forming the coarse candidates $B_{3}^{cand}$. Subsequently, the 33rd block $B_{3,3}$ refines candidates for the brute $m$-NNS (within distance threshold $\lambda$). Finally, blocks with their nearest neighbors and their observations are constructed as a contiguous arrangement for batched GPU computing.
@@
2:Output: Nearest neighbor set $J_{1:P}$
@@
6:  Update centers $C_{i}=\{\bm{c}_{p,i}\}$, where $\bm{c}_{p,i}=\frac{1}{k_{p,i}}\sum_{i=1}^{k_{p,i}}\bm{x}_{p,i}$ with $\bm{x}_{p,i}\in\bm{B}_{p,i}.$
@@
8:MPI_Allgather, gather all centers $C_{1:P}$ to each worker.
@@
19:     Calculate the distance $d(\bm{c}_{i},\bm{x}_{k})$. If $d(\bm{c}_{i},\bm{x}_{k})<\lambda$, add $\bm{x}_{k}$ to the candidate set $S_{p,i}$ for block $B_{p,i}$.
@@
23:  Searching $m$ nearest neighbors for $i$th block, i.e., $\bm{J}_{p,i}=NNS(\bm{c}_{i},S_{p,i},m)$.
@@
5.1.4. Batched Log-likelihood Computations
Once the data preprocessing is finished, we copy clusters $\bm{B}_{1:P}$, nearest neighbors $\bm{J}_{1:P}$, and their observations to GPU global memory for log-likelihood optimization. Algorithm 5 outlines the computation of a single iteration of log-likelihood, which needs to be evaluated multiple times until the prespecified optimization configuration is achieved. Here, the batched operations execute for log-likelihood computation. This involves constructing local covariance matrices $\Sigma^{lk},\Sigma^{cross}$ and $\Sigma^{con}$ using the scaled kernel $K_{\bm{\theta}}$. Subsequently, the GPU data structure is constructed in Step 4 in Figure 2. Following, batched Cholesky decompositions (POTRF) is utilized to factorize neighbor covariance matrices, $\Sigma^{con}$. Conditional mean and covariance updates are computed via batched triangular solver (TRSM) and matrix multiplications (GEMM), while batched triangular solver (TRSV) and determinant calculations derive per-block log-likelihoods $\ell_{p,i}$. Each node aggregates its local computations $\ell_{p}=\sum_{i=1}^{k_{p}}\ell_{p,i}$, and a global reduction across all nodes yields the final log-likelihood $\ell=\sum_{p=1}^{P}\ell_{p}$ using MPI_Allreduce.
Algorithm 5 Batched Block Vecchia $\ell_{llh}(B_{p},J_{p},\bm{y}_{p,B},\bm{y}_{p,J})$
1:Input: In $p$th node, $B_{p}$ cluster set, $J_{p}$ nearest neighbor set, $\bm{y}_{p,B}$ and $\bm{y}_{p,J}$ observation sets.
@@ 5.1.4. Batched Log-likelihood Computations
5.1.5. GP Prediction
During the prediction stage, the pipeline follows the estimation process, with the key difference being that in Step (2) GP calculations in Algorithm 5 are replaced by conditional simulations. Specifically, the variance vector, $\bm{\sigma}=\left(\sigma_{1},\sigma_{2},\dots,\sigma_{n^{*}}\right)^{\top}$ is subtracted from the diagonal elements of the covariance matrices $\mathbf{\Sigma}^{new}$, and $\bm{\mu}^{new}$ serves as the predicted values, $({y}_{*1},{y}_{*1},\dots,{y}_{*n^{*}})$. Next, conditional simulations are performed by drawing 1000 samples from the normal distribution $\mathcal{N}(y_{*j},\sigma_{j})$. Then, the sample mean $\tilde{\mu}_{j}$ and variance $\tilde{\sigma}_{j}^{2}$ are computed, and the 95% confidence interval is given by, $(\tilde{\mu}_{j}-z_{\alpha/2}\tilde{\sigma}_{j},\tilde{\mu}_{j}+z_{\alpha/2}\tilde{\sigma}_{j}),$ where $\alpha=0.05,j=1,2,\dots,n^{*}$, which follows the same as (Pan et al., 2025).
Overall, the workflow minimizes inter-node communication and computations by scaling and partitioning, RAC, and filtered $m$-NNS, while GPU-batched kernels exploit fine-grained parallelism for covariance and log-likelihood computations. Moreover, this algorithm incorporates anisotropic scaling for high-dimensional inputs to ensure accurate emulation for computer models. This framework enables scalable and efficient log-likelihood evaluation for large-scale GPs in the high-dimensional problem setup.
5.2. Complexity Analysis
We analyze the memory and computational complexity of SV and SBV, focusing on the estimation of GPU stage, which dominates due to repeated iterations, unlike the one-time CPU preprocessing.
@@ 6.1. Synthetic Data Simulation
Figure 4. Comparison of Vecchia-based GP methods on model fitting and prediction accuracy, i.e., Classic Vecchia (CV), Block Vecchia (BV), Scaled Vecchia (SV), Scaled Block Vecchia (SBV). Subfigures (a) and (b) evaluate the GP variants fitting (KL divergence) and prediction separately, showing that scaling and block enable better accuracy; subfigure (c) Investigate SBV with different block sizes (bs), showing that a larger $m$ and a smaller block size yield the best fitting accuracy.
The results are presented in Figure 4. In Figure 4a, the proposed SBV method achieves the lowest KL divergence, followed by SV, BV, and CV, highlighting the advantage of combining input scaling and clustering over the CV approach. The performance gains of SBV arise from its ability to address input anisotropy through joint scaling and to increase neighborhood coverage through clustering. Figure 4b shows a similar trend regarding predictive accuracy, where SBV again yields the lowest MSPE, followed by SV, BV, and CV. These results are consistent with the findings on the KL divergence and further validate the benefits of incorporating input scaling and structured clustering. In Figure 4c, we examine the impact of block size on KL divergence. Smaller blocks ($bs_{est}=12$) result in lower KL divergence compared to larger ones ($bs_{est}=50$), indicating improved approximation accuracy with finer partitioning. However, larger blocks impose heavier computational loads, which in turn allow for effective GPU utilization. Therefore, the choice of block size $bs$ introduces a trade-off between accuracy and computational cost.
@@ 6.2. Evaluation on Satellite Drag Benchmark
The satellite drag dataset in (Sun et al., 2019) is a widely used benchmark for evaluating GP-based models in high-dimensional settings (Sun et al., 2019; Katzfuss et al., 2022). It is generated from a high-fidelity simulator that models atmospheric drag coefficients in low Earth orbit (LEO). The dataset comprises 22M simulation runs for each of six primary atmospheric species: atomic oxygen ($O$), molecular oxygen ($O_{2}$), atomic nitrogen ($N$), molecular nitrogen ($N_{2}$), helium ($He$), and atomic hydrogen ($H$). The total drag coefficient is computed as a weighted average of these species. Each species-specific dataset has an 8-dimension input space: relative velocity, surface temperature, atmospheric temperature, yaw angle, pitch angle, and two accommodation coefficients.
@@
Figure 5 presents the RMSPE results for Vecchia-based GP models on the benchmark dataset. SBV consistently achieves the lowest RMSPE, demonstrating the effectiveness of combining input scaling with clustering. Increasing the number of prediction neighbors $m_{\text{pred}}$ improves accuracy; for example, SBV with $m_{\text{pred}}=400$ reduces RMSPE by approximately 15% on average compared to $m_{\text{pred}}=200$. To better understand the improved accuracy achieved by SBV, Figure 6 compares the estimated inverse lengthscales ($1/\beta_{i}$) for SV and SBV. Both methods identify the final three input dimensions as most relevant while treating the others as less influential. However, the parameter estimates produced by SBV differ significantly from those of SV, underscoring the importance of using a larger number of nearest neighbors for accurate GP fitting. From a computational and memory perspective, SV faces challenges in handling large neighbor counts due to substantial cost and memory usage increases, which limit efficient GPU utilization. In contrast, SBV remains scalable and efficient under these conditions.

### v3

@@ 5.2. Random Anchor Clustering (RAC)
4:  Randomly choose $k_{p}$ local centers $C_{p}=\{\bm{c}_{p,i}\mid i=1,2,\dots,k_{p}\}$ in $\bm{X}_{p}$, and initialize $\bm{B}_{p,j}=\{\bm{c}_{p,i}\}.$
5:  Assign $\bm{x}_{p,i}$ to $j$th cluster $\bm{B}_{p,j}$, where $j=\arg\min_{j}\|\bm{x}_{p,i}-\bm{c}_{p,j}\|^{2},j=1,2,\dots,k_{q}$.
@@
7:Return: Clusters $\bm{B}_{1:P}$
5.3. Filtered $m$-Nearest Neighbor Search (NNS)
With block sets $\bm{B}_{1:P}$, the next goal is to find the exact $m$ nearest neighbors for all block centers in high-dimensional space, where search areas are varied for each query center in the Vecchia-based GPs, see Equation (2). Existing methods (Abbasifard et al., 2014; Ferhatosmanoglu et al., 2001; Clarkson and others, 2006) for exact $m$-NNS with query-specific search areas face limitations, especially for large datasets.
To address these issues, we adopt the filtered $m$-NNS on each block $\bm{B}_{p,i}$, detailed in Algorithm 4. This approach selects a small filtered subset (a circle of radius $\lambda$) for brute-force $m$-NNS to obtain exact results while reducing computational cost. In Algorithms 1 and 4, $\mathcal{V}$ denotes the filtered $m$-NNS function, which begins by computing a Monte Carlo-based distance threshold. This threshold $\lambda$ is calculated using:
@@ 5.3. Filtered $m$-Nearest Neighbor Search (NNS)
Here, $n$ is the total number of points, $m$ is the number of nearest neighbors, $d$ is the dimensionality, and $\alpha$ is an expansion factor to account for irregularities in point distribution (e.g., $\alpha=100$ implies the candidate set is expected to be 100 times larger than $m$). Then, we prepare the small filtered subset in two steps: coarser and finer preparation. In the coarser preparation, each worker redistributes candidate clusters to the corresponding worker if the distance between two centers is within the calculated threshold using MPI_Alltoall. The complementary coarser candidate set is denoted as $\bm{B}_{p}^{cand}$ (lines 10-13 in Algorithm 4). In the finer partition, each worker independently selects finer candidates $\bm{S}_{p,i}$ for the $i$th cluster, i.e., $\bm{x}_{k}$ is added to the finer candidate set $\bm{S}_{p,i}$ if the distance between the $i$th local center and the data point $\bm{x}_{k}$ is within the threshold $\lambda$ (lines 15-22 in Algorithm 4). Finally, each local center $\bm{c}_{i}$ applies brute-force search for its $m$ nearest neighbors $\bm{J}_{p,i}$ from the finer candidates set $\bm{S}_{p,i}$. The computational cost of brute search in $\bm{S}_{p,i}$ is negligible compared to $\mathcal{O}(n)$ due to $\alpha m\ll n$ and total block count $K\ll n$, thus filtered $m$-NNS achieves computational complexity $\mathcal{O}(\alpha m+K)$ and has a simple implementation.
Figure 3 illustrates the filtered m-NNS using an example of four workers. In the coarser candidate preparation, MPI_Alltoall is utilized to distribute related blocks (within distance threshold $\lambda$) to the 33rd worker, forming the coarse candidates $\bm{B}_{3}^{cand}$. Subsequently, the 33rd block $\bm{B}_{3,3}$ refines candidates for the brute $m$-NNS (within distance threshold $\lambda$). Finally, blocks, their nearest neighbors, and their observations are arranged in a contiguous manner for batched GPU computation.
@@
2:Output: Nearest neighbor set $\bm{J}_{1:P}$
@@
6:  Update centers $\bm{C}_{i}=\{\bm{c}_{p,i}\}$, where $\bm{c}_{p,i}=\frac{1}{k_{p,i}}\sum_{i=1}^{k_{p,i}}\bm{x}_{p,i}$ with $\bm{x}_{p,i}\in\bm{B}_{p,i}.$
@@
8:MPI_Allgather, gather all centers $\bm{C}_{1:P}$ to each worker.
@@
19:     Calculate the distance $d(\bm{c}_{i},\bm{x}_{k})$. If $d(\bm{c}_{i},\bm{x}_{k})<\lambda$, add $\bm{x}_{k}$ to the candidate set $\bm{S}_{p,i}$ for block $\bm{B}_{p,i}$.
@@
23:  Searching $m$ nearest neighbors for $i$th block, i.e., $\bm{J}_{p,i}=NNS(\bm{c}_{i},\bm{S}_{p,i},m)$.
@@
5.4. Batched Log-likelihood Computations
While MPI enables scalable distributed computation across nodes, GPU acceleration delivers high single-node throughput for the SBV algorithm’s log-likelihood computations. These computationally intensive phases involve batched dense factorizations and batched matrix multiplications resulting from block conditioning. These operations are entirely offloaded to the GPU using the MAGMA batched linear algebra library. Each conditioning block in the SBV algorithm is treated as an independent batch element, which exposes fine-grained parallelism across available accelerators. Batched Cholesky factorizations and GEMM updates are executed concurrently to ensure high device occupancy. Block sizes are chosen to balance kernel launch overhead, arithmetic intensity, and memory reuse. All GPU kernels operate in double precision (FP64) to maintain numerical stability and consistency with likelihood-based Gaussian process inference. Future work will investigate the extension to lower- or mixed-precision formats.
The SBV implementation is designed to efficiently utilize the GPU memory hierarchy. After partition construction on the CPU, all associated matrices are transferred to the high-bandwidth memory (HBM) and remain resident throughout the likelihood evaluation phase. The MAGMA library is used for batched linear algebra operations (Tomov et al., 2009). Conditioning blocks are reused across multiple batched operations, thereby increasing arithmetic intensity and reducing redundant memory accesses. The batched linear algebra kernels in SBV demonstrate high arithmetic intensity as a result of repeated reuse of conditioning matrices within each partition. This characteristic makes the workload well-suited for GPU acceleration.
Host-to-device transfers are performed only during the initial data staging following partition construction. No intermediate CPU–GPU synchronization is necessary during the batched factorization and matrix update phases. Device-to-host communication is restricted to the final likelihood values and gradient quantities. Consequently, host–device transfers do not occur within the primary iterative GPU kernels and account for only a small portion of the total runtime. Maintaining data residency on the GPU throughout the computational phase minimizes PCIe/NVLink traffic. The total host-to-device transfer volume per node is proportional to the local partition size and occurs once per likelihood evaluation.
To explain this in detail, once data preprocessing is completed (Algorithm 4), the clusters $\bm{B}_{1:P}$, their corresponding nearest-neighbor sets $\bm{J}_{1:P}$, and the associated observations are transferred to the GPU global memory for subsequent log-likelihood optimization. Algorithm 5 outlines the computation of a single log-likelihood iteration, which must be evaluated multiple times until the prespecified optimization configuration is reached. Here, the batched operations execute for log-likelihood computation. This involves constructing local covariance matrices $\bm{\Sigma}^{lk},\bm{\Sigma}^{cross}$ and $\bm{\Sigma}^{con}$ using the scaled kernel $K_{\bm{\theta}}$. Subsequently, the GPU data structure is constructed in Step 4 in Figure 2. Following, batched Cholesky decompositions (POTRF) is utilized to factorize neighbor covariance matrices, $\bm{\Sigma}^{con}$. Conditional mean and covariance updates are computed via batched triangular solver (TRSM) and matrix multiplications (GEMM), while batched triangular solver (TRSV) and determinant calculations derive per-block log-likelihoods $\ell_{p,i}$. Each node aggregates its local computations $\ell_{p}=\sum_{i=1}^{k_{p}}\ell_{p,i}$, and a global reduction across all nodes yields the final log-likelihood $\ell=\sum_{p=1}^{P}\ell_{p}$ using MPI_Allreduce.
Algorithm 5 Batched Block Vecchia $\ell_{llh}(\bm{B}_{p},\bm{J}_{p},\bm{y}_{p,B},\bm{y}_{p,J})$
1:Input: In $p$th node, $\bm{B}_{p}$ cluster set, $\bm{J}_{p}$ nearest neighbor set, $\bm{y}_{p,B}$ and $\bm{y}_{p,J}$ observation sets.
@@ 5.4. Batched Log-likelihood Computations
5.5. GP Prediction
During the prediction stage, the pipeline follows the estimation process, with the key difference that, in Step (2) of Algorithm 5, GP calculations are replaced by conditional simulations. Specifically, the variance vector, $\bm{\sigma}=\left(\sigma_{1},\sigma_{2},\dots,\sigma_{n^{*}}\right)^{\top}$ is subtracted from the diagonal elements of the covariance matrices $\bm{\Sigma}^{new}$, and $\bm{\mu}^{new}$ serves as the predicted values, $({y}_{*1},{y}_{*1},\dots,{y}_{*n^{*}})$. Next, conditional simulations are performed by drawing 1000 samples from the normal distribution $\mathcal{N}(y_{*j},\sigma_{j})$. Then, the sample mean $\tilde{\mu}_{j}$ and variance $\tilde{\sigma}_{j}^{2}$ are computed, and the 95% confidence interval is given by, $(\tilde{\mu}_{j}-z_{\alpha/2}\tilde{\sigma}_{j},\tilde{\mu}_{j}+z_{\alpha/2}\tilde{\sigma}_{j}),$ where $\alpha=0.05,j=1,2,\dots,n^{*}$, which follows the same as (Pan et al., 2025).
Overall, the workflow minimizes inter-node communication and computations by scaling and partitioning, RAC, and filtered $m$-NNS, while GPU-batched kernels exploit fine-grained parallelism for covariance and log-likelihood computations. Moreover, this algorithm incorporates anisotropic scaling for high-dimensional inputs to ensure accurate emulation for computer models. This framework enables scalable and efficient evaluation of log-likelihoods for large-scale GPs in high-dimensional settings.
6. Complexity Analysis
Herein, we analyze the memory and computational complexity of SV and SBV, focusing on the GPU stage, which dominates due to repeated iterations, unlike the one-time CPU preprocessing.
@@ 7.1. Synthetic Data Simulation
(d) Relative error for RAC.
Figure 4. Comparison of Vecchia-based GP methods on model fitting and prediction accuracy, i.e., Classic Vecchia (CV), Block Vecchia (BV), Scaled Vecchia (SV), Scaled Block Vecchia (SBV). Subfigures (a) and (b) evaluate the GP variants fitting (KL divergence) and prediction separately, showing that scaling and block enable better accuracy; subfigure (c) Investigate SBV with different block sizes (bs), showing that a larger $m$ and a smaller block size yield the best fitting accuracy; subfigure (d) Clarifying that the proposed RAC has comparable and robust log-likelihood with the K-means clustering and has the advantage of computationally linear complexity.
The results are presented in Figure 4. In Figure 4a, the proposed SBV method achieves the lowest KL divergence, followed by SV, BV, and CV, highlighting the advantage of combining input scaling and clustering over the CV approach. The performance gains of SBV arise from its ability to address input anisotropy through joint scaling and to increase neighborhood coverage through clustering. Figure 4b shows a similar trend regarding predictive accuracy, where SBV again yields the lowest MSPE, followed by SV, BV, and CV. These results are consistent with the findings on the KL divergence and further validate the benefits of incorporating input scaling and structured clustering. In Figure 4c, we examine the impact of block size on KL divergence. Smaller blocks ($bs_{est}=12$) yield lower KL divergence than larger ones ($bs_{est}=50$), indicating improved approximation accuracy with finer partitioning. However, larger blocks impose heavier computational loads, thereby enabling effective GPU utilization. Therefore, the choice of block size $bs$ introduces a trade-off between accuracy and computational cost. Figure 4d illustrates the relative error and its variance of the log-likelihood computed using two clustering methods: K-means and the proposed RAC approach. To assess the robustness of RAC with respect to anchor randomness, we perform five independent random anchor selections and report both the mean error and its variance. The results show that both the approximation error and its variability decrease as the conditioning size $m$ increases, eventually becoming nearly negligible. This trend confirms the stability of RAC for sufficiently large $m$. Furthermore, larger conditioning sizes are typically required in large-scale and high-dimensional applications to maintain high approximation accuracy.
@@ 7.2. Evaluation on Satellite Drag Benchmark
The satellite drag dataset in (Sun et al., 2019) is a widely used benchmark for evaluating GP-based models in high-dimensional settings (Sun et al., 2019; Katzfuss et al., 2022). It is generated from a high-fidelity simulator that models atmospheric drag coefficients in low Earth orbit (LEO). The dataset comprises 22M simulation runs for each of six primary atmospheric species: atomic oxygen ($O$), molecular oxygen ($O_{2}$), atomic nitrogen ($N$), molecular nitrogen ($N_{2}$), helium ($He$), and atomic hydrogen ($H$). The total drag coefficient is computed as a weighted average of these species. Each species-specific dataset has an 8-dimensional input space: relative velocity, surface temperature, atmospheric temperature, yaw angle, pitch angle, and two accommodation coefficients.
@@
Figure 5 presents the RMSPE results for Vecchia-based GP models on the benchmark dataset. SBV consistently achieves the lowest RMSPE, demonstrating the effectiveness of combining input scaling with clustering. Increasing the number of prediction neighbors $m_{\text{pred}}$ improves accuracy; for example, SBV with $m_{\text{pred}}=400$ reduces RMSPE by approximately 15% on average compared to $m_{\text{pred}}=200$. To better understand the improved accuracy achieved by SBV, Figure 6 compares the estimated inverse lengthscales ($1/\beta_{i}$) for SV and SBV. Both methods identify the final three input dimensions as most relevant while treating the others as less influential. However, the parameter estimates from SBV differ significantly from those from SV, underscoring the importance of using a larger number of nearest neighbors for accurate GP fitting. From a computational and memory perspective, SV faces challenges in handling large neighbor counts because substantial increases in cost and memory usage limit efficient GPU utilization. In contrast, SBV remains scalable and efficient under these conditions.